page 57 q4 of center of mass module iam class 11 of homia22batch

page 57  q4 of center of mass module iam class 11 of homia22batch

Grade:11

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Hi,

In this question lets find out the angle of deflection (a) of two spheres from horizontal after they are being hit by another sphere.
from the geometry, sin a = r/2r, a = 30o

Applying conservation of momentum in x-direction,
mu = mv cos30 + mv cos30, v = u /\sqrt{3}

e = v cos30 / u = ½
You can now easily find out loss in kinetic energy by calculating inital and final kinetic energy in terms of u

thanks